OpenStudy (anonymous):

I need some help with finding the integral for: (x*e^x)/(x+1)^2 Does this require substitution, parts, or both?

OpenStudy (anonymous):

Possibly partial fractions

OpenStudy (anonymous):

The answer is \[e^x/(x+1)+C\] but I'm not sure how to get there

OpenStudy (anonymous):

Yeah, you can get there integration by parts.

OpenStudy (anonymous):

by parts u=xe^x, v=x+1. It's worth a try

OpenStudy (anonymous):

v=1/(x+1)

OpenStudy (dumbcow):

yes integration by parts works

OpenStudy (anonymous):

Thanks. I missed the cancellation x+1/x+1 to simplify the integral.